How do i turn 1M of hydrochloric acid to grams
1
Expert's answer
2016-01-10T02:32:15-0500
The molar mass M is a physical property defined as the mass of a given substance (chemical element or chemical compound) divided by its amount of substance. The base SI unit for molar mass is kg/mol. However, for historical reasons, molar masses are usually expressed in g/mol. The molar mass of a compound is the sum of the standard atomic weight of the atoms, which form the compound. Molar mass of HCl is:
M = MH + MCl = 1 + 35.5 = 36.5 g/mol
To turn amount of substance to grams we need multiply molar mass of substance (in mols) to amount of substance (in g/mol):
m = M * Mw = 1mol * 36.5 g/mol = 36.5 g

Answer: 36.5 g